Which of the following sets of polynomials span P2 ?

a. {t²+1, t-1, t²+t}

B. {t²+ 2, 2t²- t+1, t + 2 , t²+t+4}


1
Expert's answer
2022-03-14T18:57:16-0400

Solution (a)

 

If "\left\{ {{t^2} + 1,\,\,t - 1,\,\,{t^2} + t} \right\}\" , span P2, then "{t^2} + 1,\,\,t - 1,\,\,{t^2} + t\" must be linearly independent.

Then for, three arbitrary constants "{c_1},\,\,{c_2},\,\,{c_3}"


"\begin{array}{l}
{c_1}\left( {{t^2} + 1} \right) + {c_2}\left( {t - 1} \right) + {c_3}\left( {{t^2} + t} \right) = 0\\
{c_1}{t^2} + {c_1} + {c_2}t - {c_2} + {c_3}{t^2} + {c_3}t = 0\\
\left( {{c_1} + {c_3}} \right){t^2} + {c_2}t + {c_1} - {c_2} = 0 \cdot {t^2} + 0 \cdot t + 0
\end{array}"


This gives,


"{c_1} + {c_3} = 0\\
{c_2} = 0\\
{c_1} - {c_2} = 0"


We solve and get "{c_1} = {c_2} = {c_3} = 0\"


Therefore,  "{t^2} + 1,\,\,t - 1,\,\,{t^2} + t\"are linearly independent and hence span polynomial of degree 2. 



Solution (b)


If, "\left\{ {{t^2} + 2,\,\,\,2{t^2} - t + 1,\,\,\,t + 2,\,\,{t^2} + t + 4} \right\}\" span P2, then "{{t^2} + 2,\,\,\,2{t^2} - t + 1,\,\,\,t + 2,\,\,{t^2} + t + 4}" must be linearly independent.


Then for, four arbitrary constants "{c_1},\,\,{c_2},\,\,{c_3},\,{c_4}\"


"\begin{array}{l}
{t^2} + 2,\,\,\,2{t^2} - t + 1,\,\,\,t + 2,\,\,{t^2} + t + 4\\
{c_1}\left( {{t^2} + 2} \right) + {c_2}\left( {2{t^2} - t + 1} \right) + {c_3}\left( {t + 2} \right) + {c_4}\left( {{t^2} + t + 4} \right) = 0\\
{c_1}{t^2} + 2{c_1} + 2{c_2}{t^2} - {c_2}t + {c_2} + {c_3}t + 2{c_3} + {c_4}{t^2} + {c_4}t + 4{c_4} = 0\\
\left( {{c_1} + 2{c_2} + {c_4}} \right){t^2} + \left( { - {c_2} + {c_3} + {c_4}} \right)t + \left( {2{c_1} + {c_2} + 2{c_3} + 4{c_4}} \right) = 0 \cdot {t^2} + 0 \cdot t + 0
\end{array}"


This gives,


"\begin{array}{l}
{c_1} + 2{c_2} + {c_4} = 0 & & ... & \left( 1 \right)\\
 - {c_2} + {c_3} + {c_4} = 0 & & ... & \left( 2 \right)\\
2{c_1} + {c_2} + 2{c_3} + 4{c_4} & & ... & \left( 3 \right)
\end{array}"


These are three equations, and we have four unknowns. So, there are infinite solutions.


Therefore, the are linearly dependent and hence "{{t^2} + 2,\,\,\,2{t^2} - t + 1,\,\,\,t + 2,\,\,{t^2} + t + 4}"are linearly dependent and hence does not span polynomial of degree 2.